Putting the function into the form f(x)=a(x-h)^2+k and the function is -4x^2+32x+5
Answer by stanbon(75887)

Put the function into the form f(x)=a(x-h)^2+k
if the function is -4x^2+32x+5
----------------------
f(x) = -4x^2+32x+5
------
4x^2 - 32x - 5 = -y
------
4(x^2 - 8x + k) = -y + 5 + 4k
------
4(x^2 - 8x + 16) = -y + 5 + 4*16
------
4(x-4)^2 = -y + 69
=======
y = -4(x-4)^2 + 69
====================
Cheers,
Stan H.
===============